# LLMEval-3: A Large-Scale Longitudinal Study on Robust and Fair Evaluation of Large Language Models

### v2

@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has led to a proliferation of benchmarks designed to assess their capabilities (Chang et al. 2023; Laskar et al. 2024; Liu et al. 2025). However, these benchmarks predominantly rely on a static evaluation paradigm where models are tested on fixed, publicly available datasets (Chen, Pusarla, and Ray 2025). This approach is fundamentally vulnerable to data contamination and test set overfitting, contributing to a growing “evaluation crisis” where benchmark scores may no longer reliably reflect a model’s true, generalizable abilities (Banerjee, Agarwal, and Singh 2024; Deng et al. 2024a; Dekoninck et al. 2024). For example, GPT-4 achieved exact match rates of 52% and 57% when guessing the masked options in MMLU (Hendrycks et al. 2021a) test sets, far exceeding random chance (Deng et al. 2024b). Similarly, Qwen-1.8B has been shown to exactly replicate complete n-grams from both training and test splits of GSM8K (Cobbe et al. 2021) and MATH (Hendrycks et al. 2021b), including 25 exact 5-gram matches in the MATH test set, indicating potential undetected data leakage and memorization (Xu et al. 2024b). Furthermore, static benchmark designs, dataset contamination, and biased evaluation protocols could create misleading perceptions of LLM capabilities, undermining the reliability of current performance assessments (Banerjee, Agarwal, and Singh 2024).
@@ 2.2 Evaluation Process
  • The Outer Layer (Access Control): This layer manages authentication and authorization. We use JSON Web Tokens (JWT)(Jones, Bradley, and Sakimura 2015) to secure every API request, ensuring that only authenticated models can participate in an evaluation session. A strict Role-Based Access Control (RBAC) system prevents any cross-session or cross-user data access, isolating each evaluation.
@@ 2.3 Ranking System
To quantify answer quality, we establish a standardized scoring metric with an integer range of [0,3], ensuring consistent evaluation of response efficacy across diverse model architectures. For scoring implementation, we uniformly employ GPT-4o (OpenAI 2023) as our judge, which has demonstrated high human-machine agreement through rigorous validation (detailed in Section 4.3). This choice of a single, validated scoring model eliminates potential biases introduced by varying evaluative criteria.
@@ 3.1 Experimental Setup
Benchmark Comparison: We measure Spearman correlation between LLMEval-3 and static benchmarks (AGIEval (Zhong et al. 2024), C-Eval (Huang et al. 2023)) and perform fill-in-the-blank replay tests (1,000 questions, three attempts each) to assess contamination.
@@ 4 Related Work
Static Knowledge-Intensive Benchmarks such as MMLU (Hendrycks et al. 2021a) and C-Eval (Huang et al. 2023), provided foundational insights by assessing factual knowledge, reasoning, and cross-domain generalization in language models. These benchmarks established important baselines for measuring model performance across diverse academic domains. However, their static question sets significantly amplify the risk of data leakage, causing inflated evaluation scores and misleading perceptions of model capabilities (Banerjee, Agarwal, and Singh 2024; Xu et al. 2024a). Moreover, repeated use of fixed datasets induces benchmark-specific overfitting, with models memorizing test answers rather than achieving genuine generalization (Deng et al. 2024b). LLMEval-1 and LLMEval-2 (Zhang et al. 2024) also fall into this static category. These fundamental vulnerabilities in static evaluation have motivated researchers to explore more dynamic, robust evaluation approaches that better reflect true model generalization abilities.
Dynamic Human-Preference Evaluation such as LMSYS (Zheng et al. 2024), Chatbot Arena (Chiang et al. 2024), and AlpacaEval (Lab 2023) introduce dynamic evaluation settings by leveraging human preferences to compare model outputs in pairwise or multi-turn interactions. These approaches effectively capture conversational quality, coherence, and real-time performance, offering valuable insights beyond traditional static benchmarks. However, they often lack depth in evaluating domain-specific or complex reasoning tasks, and their reliance on non-expert crowdworkers can introduce bias and reduce reliability in assessments of specialized knowledge (Raju et al. 2024). Furthermore, the focus on conversational abilities may not adequately assess systematic knowledge coverage across academic disciplines.
LLM-as-a-Judge has emerged as a promising paradigm for scalable evaluation, where language models act as evaluators of other models’ outputs. Notable works include G-Eval, which uses GPT-4 with chain-of-thought reasoning and form-filling paradigms to assess natural language generation quality (Liu et al. 2023), and MT-Bench-101, which evaluates multi-turn dialogues through pairwise comparisons with GPT-4 justifications (Bai et al. 2024). These approaches offer significant advantages including high scalability, rapid deployment, cost-effectiveness, and demonstrated alignment with human preferences. However, most existing efforts focus primarily on eliciting better scoring mechanisms, while systematic validation and calibration of the judges, along with mitigation of inherent biases, remain underexplored (Zheng et al. 2023). Our work addresses these gaps by providing comprehensive analysis and systematic validation of the LLM-as-a-judge paradigm with rigorous human-machine agreement studies.

### v3

@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has led to a proliferation of benchmarks designed to assess their capabilities (DBLP:journals/corr/abs-2307-03109; DBLP:conf/emnlp/LaskarABRKKJB0P24; DBLP:journals/corr/abs-2506-11094). However, these benchmarks predominantly rely on a static evaluation paradigm where models are tested on fixed, publicly available datasets (DBLP:journals/corr/abs-2503-04149). This approach is fundamentally vulnerable to data contamination and test set overfitting, contributing to a growing “evaluation crisis” where benchmark scores may no longer reliably reflect a model’s true, generalizable abilities (DBLP:journals/corr/abs-2412-03597; DBLP:conf/naacl/Deng0TGC24; DBLP:journals/corr/abs-2402-02823). For example, GPT-4 achieved exact match rates of 52% and 57% when guessing the masked options in MMLU (mmlu) test sets, far exceeding random chance (leakage:deng2024). Similarly, Qwen-1.8B has been shown to exactly replicate complete n-grams from both training and test splits of GSM8K (GSM8K) and MATH (MATH), including 25 exact 5-gram matches in the MATH test set, indicating potential undetected data leakage and memorization (benchmarkleakage:xu2024). Furthermore, static benchmark designs, dataset contamination, and biased evaluation protocols could create misleading perceptions of LLM capabilities, undermining the reliability of current performance assessments (DBLP:journals/corr/abs-2412-03597).
@@ 2.2 Evaluation Process
  • The Outer Layer (Access Control): This layer manages authentication and authorization. We use JSON Web Tokens (JWT)(rfc7519) to secure every API request, ensuring that only authenticated models can participate in an evaluation session. A strict Role-Based Access Control (RBAC) system prevents any cross-session or cross-user data access, isolating each evaluation.
@@ 2.3 Ranking System
To quantify answer quality, we establish a standardized scoring metric with an integer range of [0,3], ensuring consistent evaluation of response efficacy across diverse model architectures. For scoring implementation, we uniformly employ GPT-4o (4o) as our judge, which has demonstrated high human-machine agreement through rigorous validation (detailed in Section 3.4). This choice of a single, validated scoring model eliminates potential biases introduced by varying evaluative criteria.
@@ 3.1 Experimental Setup
Benchmark Comparison: We measure Spearman correlation between LLMEval-3 and static benchmarks (AGIEval (AGIEval), C-Eval (c-eval)) and perform fill-in-the-blank replay tests (1,000 questions, three attempts each) to assess contamination.
@@ 4 Related Work
Static Knowledge-Intensive Benchmarks such as MMLU (mmlu) and C-Eval (c-eval), provided foundational insights by assessing factual knowledge, reasoning, and cross-domain generalization in language models. These benchmarks established important baselines for measuring model performance across diverse academic domains. However, their static question sets significantly amplify the risk of data leakage, causing inflated evaluation scores and misleading perceptions of model capabilities (DBLP:journals/corr/abs-2412-03597; leakage:xu2024). Moreover, repeated use of fixed datasets induces benchmark-specific overfitting, with models memorizing test answers rather than achieving genuine generalization (leakage:deng2024). LLMEval-1 and LLMEval-2 (llmeval:zhang2024) also fall into this static category. These fundamental vulnerabilities in static evaluation have motivated researchers to explore more dynamic, robust evaluation approaches that better reflect true model generalization abilities.
Dynamic Human-Preference Evaluation such as LMSYS (lmsys:zheng2024), Chatbot Arena (chatbotarena:chiang2024), and AlpacaEval (AlpacaEval) introduce dynamic evaluation settings by leveraging human preferences to compare model outputs in pairwise or multi-turn interactions. These approaches effectively capture conversational quality, coherence, and real-time performance, offering valuable insights beyond traditional static benchmarks. However, they often lack depth in evaluating domain-specific or complex reasoning tasks, and their reliance on non-expert crowdworkers can introduce bias and reduce reliability in assessments of specialized knowledge (humanprefer:raju2024). Furthermore, the focus on conversational abilities may not adequately assess systematic knowledge coverage across academic disciplines.
LLM-as-a-Judge has emerged as a promising paradigm for scalable evaluation, where language models act as evaluators of other models’ outputs. Notable works include G-Eval, which uses GPT-4 with chain-of-thought reasoning and form-filling paradigms to assess natural language generation quality (geval:liu2023), and MT-Bench-101, which evaluates multi-turn dialogues through pairwise comparisons with GPT-4 justifications (mtbench101). These approaches offer significant advantages including high scalability, rapid deployment, cost-effectiveness, and demonstrated alignment with human preferences. However, most existing efforts focus primarily on eliciting better scoring mechanisms, while systematic validation and calibration of the judges, along with mitigation of inherent biases, remain underexplored (judging:zheng2023). Our work addresses these gaps by providing comprehensive analysis and systematic validation of the LLM-as-a-judge paradigm with rigorous human-machine agreement studies.